The length of a room floor exceeds its breadth by 20 m .The area of floor remains unaltered when the length is decreased by 10 m but the breadth is increased by 5 m.The area of the floor ( in square meters)is:
Answer by TimothyLamb(4379)

---
x = length
y = width
---
x = y + 20
a = xy = (x - 10)(y + 5)
xy = xy + 5x - 10y - 50
5x - 10y = 50
---
put the system of linear equations into standard form
---
x = y + 20
5x - 10y = 50
---
x - y = 20
5x - 10y = 50
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = length = 30 m
y = width = 10 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php